I need help for my activities!! Please.Thank youd
a. -18 - [-4 - (8)]
b. -{16x - (12 - 3x)}
c. -{-9 - [-12 - (7)]} - [-15 - (-2)]
d. - {5y - [8x - (5y + 7x)]} = - {-5y - [8x - 5y -7x]}= -{5y - 8x + 5y +7x}
e. - 3(x - 2y) - (3x + 4y) - (-2x - 2y)

Answer by rothauserc(4718)

a. -18 - [-4 - (8)] =
-18 +4 + 8 =
-6
***************************************************************************
b. -{16x - (12 - 3x)} =
- ( 16x -12 + 3x ) =
-16x + 12 - 3x =
-19x + 12
***************************************************************************
c. -{-9 - [-12 - (7)]} - [-15 - (-2)] =
- ( -9 + 12 + 7 + 15 + 2 ) =
- ( 27 ) =
-27
***************************************************************************
d. - {5y - [8x - (5y + 7x)]} = - {-5y - [8x - 5y -7x]}= -{5y - 8x + 5y +7x} =
- ( 5y - ( 8x - 5y - 7x ) = - ( -5y - 8x + 5y + 7x ) = - ( 10y - x ) =
- ( 5y - 8x + 5y + 7x ) = - ( -x ) = - ( 10y - x ) =
- ( 10y - x ) = x = -10y + x =
-10y + x = x = -10y + x =
x = 0
***************************************************************************
e. - 3(x - 2y) - (3x + 4y) - (-2x - 2y) =
- 3x + 6y - 3x - 4y + 2x + 2y =
- 4x + 4y